# Low-Degree Polynomials Are Good Extractors

### v1

@@ 5.2 Low-degree polynomials extract from sumset sources
To see why this holds, consider an arbitrary fixing of $(\mathbf{U},\mathbf{U}^{\prime})=(i,j)$. Then, by an application of the data-processing inequality, it suffices to show that $(\mathbf{L}u_{i},\mathbf{L}v_{j})\approx_{m2^{-m/3}}(\mathbf{R},\mathbf{R^{\prime}})$. Towards this end, recall that for any $i,j\in[t]$, the vectors $u_{i}$ and $v_{j}$ are nonzero with disjoint supports of size $k/3$ each. Let $\mathbf{L^{\prime}}$ denote the $m\times(2m/3)$ matrix obtained by selecting columns of $\mathbf{L}$ indexed by the supports of $u_{i}$ and $v_{j}$. Then, we have that $\mathbf{L^{\prime}}\approx_{m2^{-m/3}}\mathbf{M^{\prime}}$, where $\mathbf{M^{\prime}}$ is a uniformly random $m\times(2m/3)$ matrix. To see this, note that a uniformly random vector in $\mathbb{F}_{2}^{m}$ will be linearly independent from any given collection of $32m/$ 3 vectors with probability at least $1-2^{-m/3}$, and then apply a union bound over all the $2m/3<m$ columns of $\mathbf{L^{\prime}}$. Therefore, letting $u^{\prime}_{i}$ and $v^{\prime}_{j}$ denote the restrictions of $u_{i}$ and $v_{j}$ to the coordinates in $\mathsf{supp}(u_{i})\cup\mathsf{supp}(v_{j})$, we have that $(\mathbf{L}u_{i},\mathbf{L}v_{j})\approx_{m2^{-m/3}}(\mathbf{M^{\prime}}u^{\prime}_{i},\mathbf{M^{\prime}}v^{\prime}_{j})\equiv(\mathbf{R},\mathbf{R^{\prime}})$. The last step holds because $u^{\prime}_{i}$ and $v^{\prime}_{j}$ are linearly independent, and so the random variables $\mathbf{M^{\prime}}u^{\prime}_{i}$ and $\mathbf{M^{\prime}}v^{\prime}_{j}$ are independent and uniformly distributed over $\mathbb{F}_{2}^{m}$.
@@
$\mathbf{X}^{\star}$ $\sim X^{\star}:=\{\mathsf{Ext}^{-1}_{X}(Lu_{0},a_{0}),\dots,\mathsf{Ext}^{-1}_{X}(Lu_{t},a_{t})\},$
$\mathbf{Y}^{\star}$ $\sim Y^{\star}:=\{\mathsf{Ext}^{-1}_{Y}(Lv_{0},b_{0}),\dots,\mathsf{Ext}^{-1}_{Y}(Lv_{t},b_{t})\}.$
Then, notice that the support of $\mathbf{X}^{\star}+\mathbf{Y}^{\star}$ is exactly
@@
To analyze the $\operatorname{eval}_{d}$-rank of $S^{\star}$, recall from Lemma 15 that applying linear transformations can only decrease the $\operatorname{eval}_{d}$-rank. Furthermore, note that for any $i,j\in[t]$ it holds that
@@
Furthermore, the composition $L^{-1}\circ\mathsf{Ext}$ is linear, since $L$ was linear and $\mathsf{Ext}$ comes from the leftover hash lemma. Thus,
@@
Since $|S^{\star}|\leq|\mathcal{B}_{0}|\cdot|\mathcal{B}_{1}|$, we get that $\operatorname{rank}(\operatorname{eval}_{d}(S^{\star}))=|\mathcal{B}_{0}|\cdot|\mathcal{B}_{1}|$, and so $\mathbf{X}^{\star}+\mathbf{Y}^{\star}$ has full $\operatorname{eval}_{d}$-rank. Recalling Equations 15 and 16 and the parameter settings $m=k/2$ and $\varepsilon=2^{-k/4}$, this means that the sumset source $\mathbf{X}+\mathbf{Y}$ is $\varepsilon^{\prime}$-close to a convex combination of flat sumset sources $\mathbf{X}^{\prime}+\mathbf{Y}^{\prime}$ with full $\operatorname{eval}_{d}$-rank and support sizes $|\mathsf{supp}(\mathbf{X}^{\prime})|,|\mathsf{supp}(\mathbf{Y}^{\prime})|\geq\binom{m/3}{d/2}=\binom{k/6}{d/2}$, where $\varepsilon^{\prime}=2\varepsilon+m2^{-m/3}\leq(k+2)2^{-k/12}$. ∎
Combining Lemma 17 with Lemma 18 and the triangle inequality directly yields the extractor part of Theorem 2.
5.3 Existential results for sumset extractors via a uniformly random function
In this section we prove Theorem 3, which we restate below. This resolves an open problem posed by Chattopadhyay and Liao [CL22].
See 3
In order to prove Theorem 3 we need to introduce the definition of additive energy between two sets.
Definition 15 ([TV06, Definition 2.8]).
The additive energy between two sets $\mathcal{X},\mathcal{Y}\subseteq\mathbb{F}_{2}^{n}$ is defined as
$E(\mathcal{X},\mathcal{Y}):=\sum_{w\in\mathbb{F}_{2}^{n}}|\{(x,y)\in\mathcal{X}\times\mathcal{Y}:x+y=w\}|^{2}=|\{(x,y,x^{\prime},y^{\prime})\in\mathcal{X}\times\mathcal{Y}\times\mathcal{X}\times\mathcal{Y}:x+y=x^{\prime}+y^{\prime}\}|.$
The key behind Theorem 3 is the following structural lemma, which states that every pair of independent flat sources can be written exactly as a convex combination of pairs of independent flat sources with low additive energy with slightly lower min-entropy. We prove this result by analyzing convex combinations arising from random partitions of the supports of the two independent sources. Interestingly, this high-level approach has also been exploited before in the combinatorics literature [APS17].
Lemma 19.
Let $\mathcal{X},\mathcal{Y}\subseteq\mathbb{F}_{2}^{n}$ be sets of size at least $2^{k}$ each. Then, we can partition $\mathcal{X}$ and $\mathcal{Y}$ into $T=2^{t}$ subsets $(\mathcal{X}_{i})_{i\in[T]}$ and $(\mathcal{Y}_{j})_{j\in[T]}$, respectively, such that $|\mathcal{X}_{i}|,|\mathcal{Y}_{j}|=2^{k-t}$ and $E(\mathcal{X}_{i},\mathcal{Y}_{j})\leq\ell^{2}\cdot 2^{2(k-t)}$ for all pairs $(i,j)\in[T]^{2}$, provided that $\ell\geq 4$ and $t\geq\frac{k}{\ell-1}(1+\ell/2)$.
In particular, this means that we can write $\mathbf{X}$ and $\mathbf{Y}$ uniformly distributed over $\mathcal{X}$ and $\mathcal{Y}$, respectively, as convex combinations $\mathbf{X}=\sum_{i}p_{i}\mathbf{X}_{i}$ and $\mathbf{Y}=\sum_{j}q_{j}\mathbf{Y}_{j}$ for random variables $\mathbf{X}_{i}$ and $\mathbf{Y}_{j}$ uniformly distributed over sets $\mathcal{X}_{i}$ and $\mathcal{Y}_{j}$, respectively, satisfying the properties detailed above.
@@ 5.3 Existential results for sumset extractors via a uniformly random function
Fix $\ell\geq 4$ and $t\geq\frac{k}{\ell-1}(1+\ell/2)$. Note that these constraints guarantee that $t\leq k$. Consider randomly partitioning each of $\mathcal{X}$ and $\mathcal{Y}$ (without replacement) into $T=2^{t}$ subsets $\mathcal{X}_{1},\dots,\mathcal{X}_{T}$ and $\mathcal{Y}_{1},\dots,\mathcal{Y}_{T}$, respectively, of size exactly $2^{k-t}$. Define $A_{w}=\{(x,y)\in\mathcal{X}\times\mathcal{Y}:x+y=w\}$ and $A^{i,j}_{w}=\{(x,y)\in\mathcal{X}_{i}\times\mathcal{Y}_{j}:x+y=w\}$. We show that we will sample with positive probability partitions $(\mathcal{X}_{i})$ and $(\mathcal{Y}_{j})$ such that $|A^{i,j}_{w}|\leq\ell$ for all $w\in\mathbb{F}_{2}^{n}$ and pairs $(i,j)$. In turn this implies that
$E(\mathcal{X}_{i},\mathcal{Y}_{j})=\sum_{w\in\mathbb{F}_{2}^{n}}|A_{w}^{i,j}|^{2}\leq\ell^{2}\cdot 2^{2(k-t)},$
since the set $A_{w}^{i,j}$ is non-empty for at most $|\mathcal{X}_{i}|\cdot|\mathcal{Y}_{j}|=2^{2(k-t)}$ choices of $w$, as desired.
Fix a vector $w\in\mathbb{F}_{2}^{n}$ and distinct pairs $(x^{(1)},y^{(1)}),\dots,(x^{(\ell)},y^{(\ell)})\in A_{w}$. Note that, since $x^{(i)}+y^{(i)}=x^{(j)}+y^{(j)}$, it follows that $x^{(1)},\dots,x^{(\ell)}$ are pairwise distinct, and the same holds for $y^{(1)},\dots,y^{(\ell)}$. Therefore, the probability that the pairs $(x^{(1)},y^{(1)}),\dots,(x^{(\ell)},y^{(\ell)})$ all end up in the same product set $\mathcal{X}_{i}\times\mathcal{Y}_{j}$ for some $(i,j)$ is at most [^14] This uses the fact that $\frac{a-1}{b-1}\leq\frac{a}{b}$ whenever $a\leq b$.
$\left(\frac{1}{T^{2}}\right)^{\ell-1}=2^{-2t(\ell-1)}.$
For each vector $w\in\mathbb{F}_{2}^{n}$, define $s_{w}=|A_{w}|$, and, for pairs $(x^{(a)},y^{(a)})_{a\in[\ell]}\in A_{w}^{\ell}$, let $\mathbf{Z}_{(x^{(a)},y^{(a)})_{a\in[\ell]}}$ be the indicator random variable of the event that $(x^{(a)},y^{(a)})\in\mathcal{X}_{i}\times\mathcal{Y}_{j}$ for all $a\in[\ell]$ and some indices $i,j\in[T]$. Then, taking the expectation over the sampling of the partitions $\mathcal{X}_{1},\dots,\mathcal{X}_{T}$ and $\mathcal{Y}_{1},\dots,\mathcal{Y}_{T}$, we have
$<2^{2k}\cdot 2^{k\ell}\cdot 2^{-2t(\ell-1)}$
$\leq 1.$
The second inequality uses the fact that $s_{w}\leq 2^{k}$ for all $w\in\mathbb{F}_{2}^{n}$. The third inequality holds because $\binom{2^{k}}{\ell}<2^{k\ell}$ when $\ell\geq 4$ and because $\sum_{w\in\mathbb{F}_{2}^{n}}s_{w}\leq 2^{2k}$, meaning that there are at most $2^{2k}$ vectors $w$ such that $s_{w}\neq 0$. The last inequality follows from the choice $t\geq\frac{k}{\ell-1}(1+\ell/2)$. This implies that there must exist a partition of $\mathcal{X}$ and $\mathcal{Y}$ into $\mathcal{X}_{1},\dots,\mathcal{X}_{T}$ and $\mathcal{Y}_{1},\dots,\mathcal{Y}_{T}$, respectively, such that the associated sum inside the expectation is 00, which means that $|A^{i,j}_{w}|\leq\ell$ for all $i,j\in[T]$, as desired. ∎
This lemma guarantees that to extract from any $k$-sumset source it suffices to extract from any $k$-sumset source $\mathbf{W}=\mathbf{X}+\mathbf{Y}$ such that $\mathbf{X}$ and $\mathbf{Y}$ are uniformly distributed over sets $\mathcal{X}$ and $\mathcal{Y}$, respectively, of slightly smaller size and such that the additive energy $E(\mathcal{X},\mathcal{Y})$ is appropriately small. The proof of Theorem 3 is then a consequence of Lemma 19, the following lemma of Chattopadhyay and Liao [CL22] about extraction from sumsets with low additive energy, and a union bound.
Lemma 20 ([CL22, Lemma B.2]).
For any independent flat $k$-sources $\mathbf{X},\mathbf{Y}\sim\mathbb{F}_{2}^{n}$, a uniformly random function $f:\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}$ satisfies $f(\mathbf{X}+\mathbf{Y})\approx_{\varepsilon}\mathbf{U}_{1}$ with probability at least $1-2\cdot 2^{-\frac{2\varepsilon^{2}2^{4k}}{E(\mathcal{X},\mathcal{Y})}}$.
We are now ready to prove Theorem 3.

### v2

@@ 5.2 Low-degree polynomials extract from sumset sources
To see why this holds, consider an arbitrary fixing of $(\mathbf{U},\mathbf{U}^{\prime})=(i,j)$. Then, by an application of the data-processing inequality, it suffices to show that $(\mathbf{L}u_{i},\mathbf{L}v_{j})\approx_{m2^{-m/3}}(\mathbf{R},\mathbf{R^{\prime}})$. Towards this end, recall that for any $i,j\in[t]$, the vectors $u_{i}$ and $v_{j}$ are nonzero with disjoint supports of size $m/3$ each. Let $\mathbf{L^{\prime}}$ denote the $m\times(2m/3)$ matrix obtained by selecting columns of $\mathbf{L}$ indexed by the supports of $u_{i}$ and $v_{j}$. Then, we have that $\mathbf{L^{\prime}}\approx_{m2^{-m/3}}\mathbf{M^{\prime}}$, where $\mathbf{M^{\prime}}$ is a uniformly random $m\times(2m/3)$ matrix. To see this, note that a uniformly random vector in $\mathbb{F}_{2}^{m}$ will be linearly independent from any given collection of $32m/$ 3 vectors with probability at least $1-2^{-m/3}$, and then apply a union bound over all the $2m/3<m$ columns of $\mathbf{L^{\prime}}$. Therefore, letting $u^{\prime}_{i}$ and $v^{\prime}_{j}$ denote the restrictions of $u_{i}$ and $v_{j}$ to the coordinates in $\mathsf{supp}(u_{i})\cup\mathsf{supp}(v_{j})$, we have that $(\mathbf{L}u_{i},\mathbf{L}v_{j})\approx_{m2^{-m/3}}(\mathbf{M^{\prime}}u^{\prime}_{i},\mathbf{M^{\prime}}v^{\prime}_{j})\equiv(\mathbf{R},\mathbf{R^{\prime}})$. The last step holds because $u^{\prime}_{i}$ and $v^{\prime}_{j}$ are linearly independent, so the random variables $\mathbf{M^{\prime}}u^{\prime}_{i}$ and $\mathbf{M^{\prime}}v^{\prime}_{j}$ are independent and uniformly distributed over $\mathbb{F}_{2}^{m}$.
@@
$X^{\star}$ :=$:=\{\mathsf{Ext}^{-1}_{X}(Lu_{0},a_{0}),\dots,\mathsf{Ext}^{-1}_{X}(Lu_{t},a_{t})\},$
$Y^{\star}$ :=$:=\{\mathsf{Ext}^{-1}_{Y}(Lv_{0},b_{0}),\dots,\mathsf{Ext}^{-1}_{Y}(Lv_{t},b_{t})\},$
respectively. Then, notice that the support of $\mathbf{X}^{\star}+\mathbf{Y}^{\star}$ is exactly
@@
To analyze the $\operatorname{eval}_{d}$-rank of $S^{\star}$, recall from Claim 2 that applying linear transformations can only decrease the $\operatorname{eval}_{d}$-rank. Furthermore, note that for any $i,j\in[t]$ it holds that
@@
Furthermore, the composition $L^{-1}\circ\mathsf{Ext}$ is linear, since $L$ was linear (and invertible) and $\mathsf{Ext}$ is also linear (since it comes from the leftover hash lemma). Thus,
@@
Since $|S^{\star}|\leq|\mathcal{B}_{0}|\cdot|\mathcal{B}_{1}|$, we get that $\operatorname{rank}(\operatorname{eval}_{d}(S^{\star}))=|\mathcal{B}_{0}|\cdot|\mathcal{B}_{1}|$, and so $\mathbf{X}^{\star}+\mathbf{Y}^{\star}$ has full $\operatorname{eval}_{d}$-rank. Recalling Equations 14 and 15 and the parameter settings $m=k/2$ and $\varepsilon=2^{-k/4}$, this means that the sumset source $\mathbf{X}+\mathbf{Y}$ is $\varepsilon^{\star}$-close to a convex combination of flat sumset sources $\mathbf{X}^{\star}+\mathbf{Y}^{\star}$ with full $\operatorname{eval}_{d}$-rank and support sizes $|\mathsf{supp}(\mathbf{X}^{\star})|,|\mathsf{supp}(\mathbf{Y}^{\star})|=\binom{m/3}{\lfloor d/2\rfloor}=\binom{k/6}{\lfloor d/2\rfloor}$, where $\varepsilon^{\star}=4\varepsilon+m2^{-m/3}\leq 4\cdot 2^{-k/4}+(k/2)\cdot 2^{-k/6}\leq 5k\cdot 2^{-k/6}$. This is at most $2^{-k/7}$ as long as $k$ exceeds a big enough constant $C$. Since the lemma is straightforward to obtain whenever $k\leq C$, [^16] This is because $1-2^{-2k}\leq 2^{-ck}$ whenever $k\leq C$, provided that $c>0$ is a small enough constant that depends on $C$. Thus, in this regime, the lemma amounts to simply finding a sumset source $\mathbf{W}^{\star}$ with full evald-rank that shares at least one support element with $\mathbf{W}$, which is straightforward. this completes the proof. ∎
Finally, we show how to combine Lemma 14 and Lemma 15 to get Theorem 2, restated here for convenience.
See 2
@@
Fix any $n\geq k\geq d\in\mathbb{N}$ and $\varepsilon>0$ such that $k\geq Cd(n/\varepsilon^{2})^{1/\lfloor d/2\rfloor}$, where $C$ is a sufficiently large constant, [^17] What exactly “sufficiently large” means will become clear later in the proof. and define $\varepsilon_{0}:=\varepsilon/2$. By Lemma 15, every $(n,k)$-sumset source $\mathbf{W}=\mathbf{X}+\mathbf{Y}$ is $2^{-c_{0}k}$-close to a convex combination of sumset sources $\mathbf{W}^{\star}=\mathbf{X}^{\star}+\mathbf{Y}^{\star}$ with full evald-rank and
$t:=|\mathsf{supp}(\mathbf{X}^{\star})|=|\mathsf{supp}(\mathbf{Y}^{\star})|=\binom{k/6}{\lfloor d/2\rfloor},$
where $c_{0}>0$ is some constant. By the lower bound on $k$ and the fact that $C$ is a sufficiently large constant,
$t=\binom{k/6}{\lfloor d/2\rfloor}\geq\left(\frac{k/6}{\lfloor d/2\rfloor}\right)^{\lfloor d/2\rfloor}\geq(C/3)^{\lfloor d/2\rfloor}n/\varepsilon^{2}=(C/3)^{\lfloor d/2\rfloor}n/(4\varepsilon_{0}^{2})\geq 64n/\varepsilon_{0}^{2}.$
Thus, Lemma 14 tells us that a random degree $d$ polynomial $f:\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}$ is an $(\varepsilon_{0}+2^{-c_{0}k})$-extractor for $(n,k)$-sumset sources with $k\geq Cd(n/\varepsilon^{2})^{1/\lfloor d/2\rfloor}$, except with probability at most $2\cdot 2^{-\varepsilon_{0}^{2}t^{2}/32}$.
We now argue that $\varepsilon_{0}+2^{-c_{0}k}\leq\varepsilon$. For this, it suffices to argue that $2^{-c_{0}k}\leq\varepsilon_{0}$ (which was set to $\varepsilon/2$). Thus, given the lower bound on $k$, we just need $2^{-c_{0}Cd(n/\varepsilon^{2})^{1/\lfloor d/2\rfloor}}\leq\varepsilon$. For this, it suffices to show $2^{-c_{0}Cd(1/\varepsilon)^{1/\lfloor d/2\rfloor}}\leq\varepsilon$. And if $C$ is a sufficiently large constant $C\geq(\log_{2}e)/(2c_{0})$, then it suffices to show $2^{-(\log_{2}e)(d/2)(1/\varepsilon)^{1/\lfloor d/2\rfloor}}\leq\varepsilon$, or $e^{\lfloor d/2\rfloor(1/\varepsilon)^{1/\lfloor d/2\rfloor}}\geq 1/\varepsilon$, which is $e^{(1/\varepsilon)^{1/\lfloor d/2\rfloor}}\geq(1/\varepsilon)^{1/\lfloor d/2\rfloor}$. And this follows from the standard inequality $x\leq e^{x}$ for all real $x$.
Finally, it is easy to verify that $2\cdot 2^{-\varepsilon_{0}^{2}t^{2}/32}\leq 2^{-\varepsilon^{2}\binom{k/C}{2\lfloor d/2\rfloor}}$, using our definitions of $\varepsilon_{0}$ and $t$ above, and the fact that $k\geq Cd(n/\varepsilon^{2})^{1/\lfloor d/2\rfloor}$ for a sufficiently large constant $C$. ∎
5.3 Impossibility results
Above, we proved that random low-degree polynomials are good extractors for sumset sources (Theorem 2). Given this result, it is natural to ask just how tight it is: in particular, can one show that every low-degree polynomial is nevertheless constant on some relatively large sumset source? As it turns out, the work of Cohen and Tal [CT15] already gives an answer to this question. Indeed, they show that every low-degree polynomial is constant on some relatively large affine source, which is a special case of a sumset source. Thus, their result already implies that every degree $d$ polynomial $f:\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}$ is constant on some $(n,k)$-sumset source $\mathbf{W}$ with $k\geq\Omega(dn^{\frac{1}{d-1}})$, proving that Theorem 2 is tight up to the constant $2$ in the exponent.
In this section, we instead ask whether such a result can be obtained for two independent sources – which can be viewed as the other well-known specialization of sumset sources. We prove such a result below, and show that it holds even if we consider the following (more general) family of functions $f:\mathbb{F}_{2}^{n}\times\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}$.
Definition 15 (Left-degree).
We say that a function $f:\mathbb{F}_{2}^{n}\times\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}$ has left-degree $d$ if for any fixed $y\in\mathbb{F}_{2}^{n}$, the restricted function $f(\cdot,y)$ can be written as an $\mathbb{F}_{2}$-polynomial of degree $\leq d$.
Note that every polynomial of degree $\leq d$ also has left degree $\leq d$, but a function with left degree $\leq d$ can be much more complex than a degree $\leq d$ polynomial. Now, given this definition, we prove the following impossibility result for two-source dispersers that are computed by functions with low left-degree.
Theorem 7.
There is a constant $c>0$ such that for every positive integers $n$ and $d\leq c\log n$, the following holds. For any function $f:\mathbb{F}_{2}^{n}\times\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}$ with left-degree $\leq d$, there exist independent $(n,k)$-sources $\mathbf{X},\mathbf{Y}$ with $k\geq cdn^{1/d}$ such that $f(\mathbf{X},\mathbf{Y})$ is constant.
Before we prove Theorem 7, we present a corollary for degree $\leq d$ polynomials.
Corollary 6.
There is a constant $c>0$ such that for every positive integers $n$ and $d\leq c\log n$, the following holds. For every polynomial $f:\mathbb{F}_{2}^{2n}\to\mathbb{F}_{2}$ with degree $\leq d$, there exist independent $(n,k)$-sources $\mathbf{X},\mathbf{Y}$ with $k\geq cdn^{\frac{1}{d-1}}$ such that $f(\mathbf{X},\mathbf{Y})$ is constant.
Proof.
Every polynomial $f:\mathbb{F}_{2}^{2n}\to\mathbb{F}_{2}$ of degree at most $d$ can be written as
$f(x,y)=g(x,y)+h(x),$
such that $g$ has degree at most $d$ and does not feature monomials of $f$ containing only variables from $\{x_{1},\dots,x_{n}\}$, and $h$ collects all remaining monomials. Note that $g$ has left-degree at most $d-1$. Thus, by Theorem 7, there exist independent $(n,k)$-sources $\mathbf{X},\mathbf{Y}$ with $k\geq cdn^{\frac{1}{d-1}}$ such that $g(\mathbf{X},\mathbf{Y})$ is constant.
Next, without loss of generality, assume that $\Pr[h(\mathbf{X})=0]\geq 1/2$. In this case, the random variable $\mathbf{X}^{\prime}=(\mathbf{X}\mid h(\mathbf{X})=0)$ satisfies
$H_{\infty}(\mathbf{X}^{\prime})\geq H_{\infty}(\mathbf{X})-1\geq k-1\geq c^{\prime}dn^{\frac{1}{d-1}}$
for some absolute constant $c^{\prime}>0$. Since $f(\mathbf{X}^{\prime},\mathbf{Y})=g(\mathbf{X}^{\prime},\mathbf{Y})$ is constant, the desired result follows. ∎
Now, in order to prove Theorem 7, we revisit the argument used by Cohen and Tal [CT15] in the context of affine sources. In particular, we will use the following result from [CT15, Appendix B], which is a consequence of the Chevalley-Warning theorem. It says that if a low-degree polynomial $f$ vanishes on a subspace $V$, then it also vanishes on many shifts of that subspace.
Lemma 16 ([CT15, Appendix B]).
Consider a subspace $V\subseteq\mathbb{F}_{2}^{n}$ of dimension $t$ and a polynomial $f:\mathbb{F}_{2}^{n}\to\mathbb{F}_{2}$ of degree at most $d$ such that $f(v)=0$ for all $v\in V$. Then, it holds that
|$\left|\{x\in\mathbb{F}_{2}^{n}:\forall v\in V,f(x+v)=0\}\right|\geq 2^{n-\sum_{j=0}^{d-1}(d-j)\binom{t}{j}}.$
With this tool in hand, we are ready to prove Theorem 7.